Dip update 80/n
[Orig: Sept 11, 2017]
LCO update: Bad weather again at OGG, so only TFN to report.
The latest TFN measurement shows the flux is up a bit from the night before - about at the same level as it was two nights ago. It will be interesting to see if the event will turn out to be symmetric (i.e. rise back up to normal the same way it fell), or if it will show more variation on its way to recovery. Only time will tell!
Note: I have turned off the detrending I had previously imposed on the light curve extraction. As mentioned in a previous post, this is a simple method (perhaps too simple), and we are working on a better technique. After some thought, I think its best to show the non-detrended data because I have learned that this simple detrending model applies slightly different shifts to the data every time the data are analyzed and posted on this blog (NB: all 'shifts' are small, within the error bars). This lack of consistency is not the way I would like to represent the data as new updates get posted on this site. Once we have made progress on properly detrending, we will be sure to update the community here with details.
~Tabby and team
PS: These observations are happening because of the wonderful backers of our 2016 Kickstarter project. The Kickstarter campaign has ended, but we are still accepting donations to purchase additional observing time on the LCO 0.4m network. Thanks in advance for your support!